# Marginally interpretable spatial logistic regression with bridge processes

### v2

@@ 4.1 Comparison with existing approaches
SpGEE, spatial generalized estimating equation method of Cattelan and Varin, (2018); BrP, bridge process random effect model; GP, Gaussian process random effect model; RMSE, root mean squared error.
The simulation results are summarized in Table 1 in terms of population-averaged estimate $\hat{\beta}^{M}=(\hat{\beta}^{M}_{0},\hat{\beta}^{M}_{1})^{\mathrm{\scriptscriptstyle T}}$ and site-specific estimate $\hat{\beta}=(\hat{\beta}_{0},\hat{\beta}_{1})^{\mathrm{\scriptscriptstyle T}}$, where posterior mean estimates are used for Bayesian approaches. All methods yield consistent estimates in general, but exhibit some differences in terms of mean square error and coverage probabilities depending on underlying data-generating models.
Under the bridge process data-generating model, the spatial generalized estimating equation method gives generally higher mean squared error and narrower confidence intervals of $\hat{\beta}^{M}$, leading to low coverage probability, especially under 50% for $\hat{\beta}^{M}_{1}$. The intercept coverage probability of the spatial generalized estimating equation is slightly lower than the nominal level as the spatial dependence increases, which is consistent with the findings of Cattelan and Varin, (2018). Here the robust standard error correction cannot be easily applied due to the lack of independent replicates; see Cattelan and Varin, (2018) for a discussion. The proposed bridge process random effect model performs well for both $\hat{\beta}^{M}$ and $\hat{\beta}$ as the model is correctly specified, and the Gaussian process random effect model gives a similar site-specific estimate to the proposed model.
Under the Gaussian copula process data-generating model, the spatial generalized estimating equation method provides reasonable coverage probabilities compared to the proposed model, yet has a slightly higher mean square error. The proposed model gives a wider credible interval for $\hat{\beta}^{M}_{1}$, and a similar pattern can be observed for $\hat{\beta}_{1}$ but to a milder degree. In all settings, the mean square error of estimators increases as spatial dependence becomes stronger.
@@
Data generation SpGEE BrP GP
Bridge process $\rho=0.05$ Test log-likelihood ×$\times 10^{2}$63.446 -53.073 -53.069
Test AUC ×$\times 10^{2}$ 68.304 80.130 80.137
Bridge process $\rho=0.1$ Test log-likelihood ×$\times 10^{2}$62.390 -49.963 -49.963
Test AUC ×$\times 10^{2}$ 68.541 81.931 81.932
Gaussian copula $\rho=0.05$ Test log-likelihood ×$\times 10^{2}$63.545 -52.564 -52.572
Test AUC ×$\times 10^{2}$ 68.280 80.866 80.861
Gaussian copula $\rho=0.1$ Test log-likelihood ×$\times 10^{2}$62.561 -49.108 -49.103
Test AUC ×$\times 10^{2}$ 68.520 83.033 83.037
Test AUC, area under the receiver operating characteristic curve based on test dataset; SpGEE, spatial generalized estimating equation method of Cattelan and Varin, (2018); BrP, bridge process random effect model; GP, Gaussian process random effect model. The Monte Carlo standard errors are all less than 0.5 for $\rho=0.05$ and all less than 0.7 for $\rho=0.1$.
Table 2 summarizes predictive performance based on held-out data of size $n_{\mathrm{test}}=50$, $N_{\mathrm{test}}=500$. We compare test log-likelihood averaged over $N_{\mathrm{test}}=500$ binary data and area under the receiver operating characteristic curve based on test data (test AUC). For the spatial generalized estimating equation method, we used $\mathrm{logit}^{-1}(x_{ij}^{{\mathrm{\scriptscriptstyle T}}}\hat{\beta}^{M})$ as predicted probabilities and for Bayesian approaches, we used the posterior predictive mean. The results suggest that the bridge process and Gaussian process random effect models produce virtually identical predictions which are superior to marginal methods that cannot yield site-specific predictions.
@@ 4.2 Scalability analysis with low-rank dependence structure
Next, we conduct another simulation study to analyze the benefits of scalable computing strategies described in Section 3.2 when the number of spatial locations $n$ becomes large. In the random effects model of the bridge process, we compare how the choice of the correlation kernel $\mathcal{R}$ affects the inference, prediction and sampling efficiency of the parameters.
@@
We generate data based on the bridge process random effects model (4). We choose spatial locations uniformly at random from the square domain $\mathscr{S}=[0,2]^{2}$ to decide training and test locations with sizes $(n_{\mathrm{train}},n_{\mathrm{test}})\in\{(200,50),(800,200)\}$. Similarly to previous simulation settings, we set $\phi=0.7$ and use the Matérn correlation kernel $\mathcal{R}_{M}(s,s^{\prime})=(1+\|s-s^{\prime}\|_{2}/\rho)\exp(-\|s-s^{\prime}\|_{2}/\rho)$ with known smoothness $1.5$ and unknown range parameter $\rho$, where the true $\rho$ are set as $\rho=0.05$ or $\rho=0.1$. We set $N_{i}=10$ for all locations, set $p=2$ including the intercept, with $x_{ij}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}N(0,1)$. The true fixed effect coefficient is set as $\beta=(\beta_{0},\beta_{1})^{\mathrm{\scriptscriptstyle T}}=(0,1)^{\mathrm{\scriptscriptstyle T}}$, and the corresponding true population-averaged coefficient is $\beta^{M}=(\beta_{0}^{M},\beta_{1}^{M})^{\mathrm{\scriptscriptstyle T}}=(0,0.7)^{\mathrm{\scriptscriptstyle T}}$. Denoting $N=\sum_{i=1}^{n}N_{i}$, we have training data sizes $(n,N)\in\{(200,2000),(800,8000)\}$ with two different ranges $\rho\in\{0.05,0.1\}$, and this data generation process is repeated 200 times.
@@
We consider 3 different correlation kernels, one with a full-rank Matern kernel $\mathcal{R}_{M}(s,s^{\prime})=(1+\|s-s^{\prime}\|_{2}/\rho)\exp(-\|s-s^{\prime}\|_{2}/\rho)$ and two with a low-rank kernel (8) with $q=49$ and $q=100$. Within the domain $[0,2]^{2}$, the knot locations $\{\tilde{s}_{k}\}_{k=1}^{q}$ are selected as $\{0.1,0.4,\dots,1.6,1.9\}^{2}$ for $q=49$ and $\{0.1,0.3,\dots,1.7,1.9\}^{2}$ for $q=100$, respectively. The range parameter is assumed to be unknown and we adopt the same prior specification as in Section 4.1. For the Markov chain Monte Carlo, we run 6,000 iterations, with the first 1,000 samples discarded as burn-in and 5,000 samples are saved without thinning.
@@
Figure 2: Boxplot summaries of scalability analysis simulation result for $\rho=0.05$ (moderate spatial dependence) based on 200 replicated datasets.
@@
Figure 3: Boxplot summaries of scalability analysis simulation result for $\rho=0.1$ (strong spatial dependence) based on 200 replicated datasets.
@@
The simulation results are summarized in Figure 2 and Figure 3. In terms of predictive performance based on average test log-likelihood and test AUC, the full-rank correlation kernel performs better than the low-rank ones for $\rho=0.05$ but similarly for $\rho=0.1$ and $q=100$. This is an expected result, since low-rank methods typically yield predictions that smooth over small-scale patterns (Datta et al.,, 2016), thus the difference is more emphasized when $\rho=0.05$. We also compare the sampling efficiency of $\beta^{M}$ and $\beta$ in terms of multivariate effective sample size (Vats et al.,, 2019) divided by running time (mESS/sec). As the number of spatial locations increases from $n=200$ to $n=800$, the mESS/sec between full-rank and low-rank differs in an order of magnitude for both $\beta^{M}$ and $\beta$, clearly showing the computational benefits of the low-rank correlation kernel. The difference is mainly attributed to the running time difference, as mESS values not standardized by time are similar.
@@ 5 Childhood malaria prevalence in Gambia
We illustrate the proposed model by analyzing malaria data among children in Gambia (Thomson et al.,, 1999), The dataset, available in R package geoR (Diggle et al.,, 1998), contains $\sum_{i=1}^{n}N_{i}=2035$ children’s malaria infection statuses from $n=65$ villages, along with covariates including age, bed net use, net treatment with insecticides, a satellite-derived measure of greenness, and the presence of a health center in the villages. The study of population-averaged associations between malaria prevalence and covariates is useful for nationwide decision-making; for example, on supplying bed nets. The prediction of malaria prevalence in new locations is also of substantial interest. The previous literature suggests residual spatial dependence in these data (Diggle et al.,, 2002; Bai et al.,, 2014; Cattelan and Varin,, 2018).
We choose the same set of variables as Cattelan and Varin, (2018), reproducing their results based on spatial generalized estimating equations, and comparing with bridge and Gaussian process random effect logistic models with an exponential correlation kernel. We use the same priors as in the simulation study (Gelman,, 2006; Gelman et al.,, 2008) except for the range parameter $\rho$. For the prior on the spatial range $\rho$ in the exponential correlation kernel $\mathcal{R}(s,s^{\prime})=\exp(\|s-s^{\prime}\|_{2}/\rho)$, we assign a uniform prior $\rho\sim\mathrm{Unif}(0.01,100)$ for the bridge and Gaussian process random effects models. We run three Markov chains with 11,000 iterations, where the first 1,000 samples are discarded as burn-in and 10,000 samples are saved without thinning. The running time based on a pure R implementation is about 2 mins for the bridge process random effect model and about 1.5 mins for the Gaussian process random effect model under the Apple M1 3.20GHz CPU environment.
@@
Population-average estimate Site-specific estimate
Variable Spatial GEE Bridge process RE Bridge process RE GP RE
Intercept 6.95 (0.39, 13.52) 1.76 (-2.68, 6.08) 2.39 (-3.47, 7.92) 2.43 (-3.07, 7.84)
Age (years) 0.22 (0.14, 0.30) 0.18 (0.10, 0.28) 0.24 (0.16, 0.33) 0.24 (0.16, 0.33)
Net-use -0.39 (-0.66, -0.11) -0.27 (-0.55, -0.04) -0.36 (-0.67, -0.05) -0.36 (-0.67, -0.06)
Treated net -0.32 (-0.65, 0.01) -0.27 (-0.61, 0.02) -0.36 (-0.75, 0.03) -0.36 (-0.75, 0.03)
Green -0.36 (-0.63, -0.08) -0.10 (-0.27, 0.07) -0.13 (-0.34, 0.09) -0.13 (-0.34, 0.08)
@@
Health center -0.26 (-0.60, 0.08) -0.22 (-0.59, 0.11) -0.30 (-0.74, 0.15) -0.30 (-0.71, 0.12)
Spatial GEE, spatial generalized estimating equation method of Cattelan and Varin, (2018); RE, random effects; GP, Gaussian process.
The results are summarized in Table 3. Trace plots of the bridge process and Gaussian process random effect models are shown in Figure A.4, both showing excellent convergence. The site-specific estimates using the bridge and Gaussian process random effect models give almost identical results. From the bridge process model, the posterior mean of the attenuation parameter is $\hat{\phi}=0.75$ with 95% credible interval $(0.46,0.97)$. The spatial generalized estimating equation and the proposed method give similar results for population-averaged estimates, except for the intercept and greenness. Specifically, the significance of the intercept and greenness coefficients at the 0.05 level differs across two methods. This discrepancy can be partly explained by simulation results, where the coverage of the intercept estimate $\hat{\beta}^{M}_{0}$ under the true $\beta_{0}^{M}=0$ is lower than the nominal level with the spatial generalized estimating equation method but not with the proposed method.
In addition to the estimated population-averaged and site-specific coefficients, we have posterior mean $\hat{\rho}=33.3(14.2,47.2)$ and posterior mode $\hat{\rho}^{\mathrm{mode}}=13.5$ for the bridge process, and posterior mean $\hat{\rho}=29.9(13.4,40.1)$ and posterior mode $\hat{\rho}^{\mathrm{mode}}=13.5$ for the Gaussian process random effect models, where parentheses correspond to the posterior interquartile range. The widely applicable information criterion (Gelman et al.,, 2014), conditional on random effects, gives 2329.5 (standard error 40.2) for the bridge process random effect model and 2326.3 (standard error 40.0) for the Gaussian process random effect model. Although the Gaussian process model shows slightly better predictive performance, the difference is negligible when considering the benefit of dual interpretation offered by the bridge process model.

### v3

@@ 4.1 Comparison with existing approaches
SpGEE, spatial generalized estimating equation method of Cattelan and Varin, (2018); BrP, proposed bridge process model; BrP-FB, bridge process model with fully Bayesian approach on $\phi$.
Table 2: Comparison of estimated site-specific effects based on 200 simulations in terms of bias, root mean squared error (RMSE), average length of 95% confidence or credible interval (CI.95), and coverage probabilities at nominal level 0.95 (Cover).
Site-specific ($\hat{\beta}_{1}$) Data from bridge process Data from Gaussian copula
Setting Method Bias RMSE CI.95 Cover Bias RMSE CI.95 Cover
$\phi=0.7$, $\rho=0.05$ GP 0.002 0.068 0.275 95.5% -0.001 0.063 0.270 98.5%
BrP 0.003 0.068 0.275 96.5% -0.001 0.063 0.270 98.5%
BrP-FB 0.004 0.068 0.281 96.5% 0.001 0.063 0.275 98.5%
$\phi=0.7$, $\rho=0.1$ GP -0.001 0.074 0.271 93.5% -0.002 0.065 0.267 97.0%
BrP -0.002 0.074 0.271 94.0% -0.001 0.066 0.266 97.5%
BrP-FB 0.001 0.074 0.279 94.5% -0.002 0.065 0.276 97.0%
$\phi=0.9$, $\rho=0.05$ GP 0.001 0.064 0.251 96.0% -0.003 0.060 0.243 95.0%
BrP 0.000 0.063 0.250 95.5% -0.005 0.061 0.242 95.5%
BrP (FB) 0.001 0.063 0.252 96.0% -0.003 0.060 0.245 96.0%
$\phi=0.9$, $\rho=0.1$ GP 0.000 0.066 0.249 94.5% -0.001 0.059 0.241 96.0%
BrP -0.001 0.066 0.248 94.5% -0.003 0.059 0.241 95.0%
BrP-FB 0.001 0.066 0.254 94.5% 0.000 0.059 0.246 96.5%
GP, Gaussian process random effect model; BrP, proposed bridge process model; BrP-FB, bridge process model with fully Bayesian approach on $\phi$.
The simulation results are summarized in terms of the estimated population-averaged effects $\hat{\beta}_{1}^{\textsc{m}}$ in Table 1 and the site-specific effects $\hat{\beta}_{1}$ in Table 2. For summaries of the intercept terms, see Supplementary Section A.4. We first remark that when the data are generated from bridge processes, the root mean squared error (RMSE) of $\hat{\beta}_{1}^{\textsc{m}}$ in Table 1 is substantially larger than that under the Gaussian copula, whereas the RMSEs of the conditional effect $\hat{\beta}_{1}$ in Table 2 are comparable across the two data generating mechanisms. This suggests that there is a large variability of population-averaged effect $\hat{\beta}_{1}^{\textsc{m}}$ under the bridge process data-generating scheme, compared to the Gaussian copula data-generating scheme.
In terms of the estimated population-averaged effects $\hat{\beta}_{1}^{\textsc{m}}$, the SpGEE and BrP produce generally similar results, whereas BrP-FB shows substantial differences. Notably, BrP-FB exhibits substantial bias when $\phi=0.9$ for both data generation scenarios, and when $\phi=0.7$, the bias becomes confounded with the influence of the prior $p(\phi)$. In contrast, both SpGEE and BrP show relatively robust results across different $(\phi,\rho)$ and data settings, highlighting the advantage of the empirical Bayes approach to find $\hat{\phi}$ based on composite likelihood. Regarding confidence and credible interval, BrP-FB produces much wider intervals than both SpGEE and BrP. Although BrP-FB yields the desired coverage when the data are generated from a bridge process random effects, it also results in 100% coverage under Gaussian copula data, suggesting that uncertainty quantification with BrP-FB is highly sensitive to model misspecification.
For the estimated site-specific effects $\hat{\beta}_{1}$, all methods GP, BrP, and BrP-FB lead to highly similar results. While there are slight overcoverages under the Gaussian copula data generation setting with $\phi=0.7$, generally there are almost no biases and all three methods achieve the desired coverage level. Table 3 summarizes predictive performance based on held-out data of size $n_{\mathrm{test}}=50$, $N_{\mathrm{test}}=500$. Across three methods GP, BrP, and BrP-FB, we compare test log-likelihood averaged over $N_{\mathrm{test}}=500$ binary data and area under the receiver operating characteristic curve based on test data (test AUC) using the posterior predictive mean. The results suggest that the GP, BrP, and BrP-FB models produce virtually identical predictions. In summary, the results demonstrate that BrP with an empirical Bayes estimate of $\phi$ provides a marginal effect estimate comparable to SpGEE, and at the same time yield site-specific estimates and predictions that are virtually identical to the Gaussian process random effects model.
@@
Data from bridge process Data from Gaussian copula
Setting Method Test loglik $\times 10^{2}$ Test AUC $\times 10^{2}$ Test loglik $\times 10^{2}$ Test AUC $\times 10^{2}$
$\phi=0.7$, $\rho=0.05$ GP -53.069 80.137 -52.572 80.861
BrP -53.076 80.129 -52.571 80.863
BrP-FB -53.073 80.130 -52.564 80.866
$\phi=0.7$, $\rho=0.1$ GP -49.963 81.932 -49.103 83.037
BrP -49.964 81.929 -49.108 83.034
BrP-FB -49.963 81.931 -49.108 83.033
$\phi=0.9$, $\rho=0.05$ GP -57.410 76.437 -58.680 75.339
BrP -57.415 76.433 -58.680 75.339
BrP-FB -57.408 76.440 -58.671 75.351
$\phi=0.9$, $\rho=0.1$ GP -55.445 77.664 -57.224 76.567
BrP -55.445 77.667 -57.229 76.563
BrP-FB -55.446 77.663 -57.223 76.566
GP, Gaussian process random effect model; BrP, proposed bridge process model; BrP-FB, bridge process model with fully Bayesian approach on $\phi$. The Monte Carlo standard errors of Test loglik $\times 10^{2}$ and Test AUC $\times 10^{2}$ are all less than 0.5 for $\rho=0.05$ and all less than 0.7 for $\rho=0.1$.
@@ 4.2 Scalability analysis with low-rank dependence structure
Next, we conduct another simulation study to analyze the benefits of scalable computing strategies described in Section 3.3 when the number of spatial locations $n$ becomes large. Based on the proposed bridge process random effects model, we compare how the low-rank structure of the correlation kernel affects the prediction and sampling efficiency of the parameters.
@@
We generate data based on the bridge process random effects model (6). We choose spatial locations uniformly at random from the square domain $\mathscr{S}=[0,2]^{2}$ to decide training and test locations with sizes $(n_{\mathrm{train}},n_{\mathrm{test}})\in\{(200,50),(800,200)\}$. Similarly to previous simulation settings, we set $\phi=0.7$ and use the Matérn correlation kernel $\mathcal{R}_{M}(s,s^{\prime})=(1+\|s-s^{\prime}\|_{2}/\rho)\exp(-\|s-s^{\prime}\|_{2}/\rho)$ with known smoothness $1.5$ and unknown range parameter $\rho$, where the true $\rho$ are set as $\rho=0.05$ or $\rho=0.1$. We set $N_{i}=10$ for all locations, set $p=2$ including the intercept, with $x_{ij}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}N(0,1)$. The true fixed effect coefficient is set as $\beta=(\beta_{0},\beta_{1})^{\mathrm{\scriptscriptstyle T}}=(0,1)^{\mathrm{\scriptscriptstyle T}}$. Denoting $N=\sum_{i=1}^{n}N_{i}$, we have training data sizes $(n,N)\in\{(200,2000),(800,8000)\}$ with two different ranges $\rho\in\{0.05,0.1\}$, and this data generation process is repeated 200 times.
@@
We consider BrP (empirical Bayes) and BrP-FB with 3 different correlation kernels, one with a full-rank Matern kernel $\mathcal{R}_{M}(s,s^{\prime})=(1+\|s-s^{\prime}\|_{2}/\rho)\exp(-\|s-s^{\prime}\|_{2}/\rho)$ and two with a low-rank kernel (12) with $q=49$ and $q=100$. Within the domain $0,2]^{2}$, the knot locations $\{\tilde{s}_{k}\}_{k=1}^{q}$ are selected as $\{0.1,0.4,\dots,1.6,1.9\}^{2}$ for $q=49$ and $\{0.1,0.3,\dots,1.7,1.9\}^{2}$ for $q=100$, respectively. The range parameter is assumed to be unknown and we adopt the same prior specification as in Section 4.1. For the Markov chain Monte Carlo, we run 6,000 iterations, with the first 1,000 samples discarded as burn-in and 5,000 samples are saved without thinning.
@@
Figure 2: Boxplot summaries of scalability analysis simulation result for $\rho=0.05$ (moderate spatial dependence) based on 200 replicated datasets. The sampling efficiency is displayed in log scale.
@@
Figure 3: Boxplot summaries of scalability analysis simulation result for $\rho=0.1$ (strong spatial dependence) based on 200 replicated datasets. The sampling efficiency is displayed in log scale.
@@
The simulation results for BrP-FB are summarized in Figure 2 and Figure 3. In terms of predictive performance based on average test log-likelihood and test AUC, the full-rank correlation kernel performs better than the low-rank ones for $\rho=0.05$ but similarly for $\rho=0.1$ and $q=100$. This is an expected result, since low-rank methods typically yield predictions that smooth over small-scale patterns (Datta et al., , 2016), thus the difference is more emphasized when $\rho=0.05$. We also compare the sampling efficiency of $\bm{\beta}^{\textsc{m}}$} and $\bm{\beta}$ in terms of multivariate effective sample size (Vats et al., , 2019) divided by running time (mESS/sec). As the number of spatial locations increases from $n=200$ to $n=800$, the mESS/sec between full-rank and low-rank differs in an order of magnitude for both $\bm{\beta}^{\textsc{m}}$} and $\bm{\beta}$, clearly showing the computational benefits of the low-rank correlation kernel. The difference is mainly attributed to the running time difference, as mESS values not standardized by time are similar. As Table 3 suggests, the predictive performance under BrP is virtually identical to that under BrP-FB, and the sampling efficiency results for BrP (both $\bm{\beta}^{\textsc{m}}$} and $\bm{\beta}$) are highly similar to those for $\bm{\beta}$ under BrP-FB, and are omitted for better clarity.
@@ 5 Childhood malaria prevalence in Gambia
We illustrate the proposed model by analyzing malaria data among children in Gambia (Thomson et al., , 1999), The dataset, available in R package geoR (Diggle et al., , 1998), contains $\sum_{i=1}^{n}N_{i}=2035$ children’s malaria infection statuses from $n=65$ villages, along with covariates including age, bed net use, net treatment with insecticides, a satellite-derived measure of greenness, and the presence of a health center in the villages. The study of population-averaged associations between malaria prevalence and covariates is useful for nationwide decision-making; for example, on supplying bed nets. The prediction of malaria prevalence in new locations is also of substantial interest. The previous literature suggests residual spatial dependence in these data (Diggle et al., , 2002; Bai et al., , 2014; Cattelan and Varin, , 2018). In Supplementary Section A.5, we provide a more detailed investigation of residual spatial dependence.
We choose the same set of variables as Cattelan and Varin, (2018), reproducing their results based on spatial generalized estimating equations, and comparing with bridge process (empirical Bayes for $\phi$) and Gaussian process random effect logistic models with an exponential correlation kernel. We use the same priors as in the simulation study (Gelman, , 2006; Gelman et al., , 2008) except for the range parameter $\rho$. For the prior on the spatial range $\rho$ in the exponential correlation kernel $\mathcal{R}_{E}(s,s^{\prime})=\exp(\|s-s^{\prime}\|_{2}/\rho)$, we assign a uniform prior $\rho\sim\mathrm{Unif}(0.01,100)$ for the bridge and Gaussian process random effects models. We run three Markov chains with 11,000 iterations, where the first 1,000 samples are discarded as burn-in and 10,000 samples are saved without thinning. The wall-clock running time is about 2 mins for the bridge process random effect model with empirical Bayes approach, including the time for finding $\hat{\phi}$ which took approximately 1 min, and about 1.5 mins for the fully Bayesian Gaussian process random effect model under the Apple M1 3.20GHz CPU environment.
@@
Population-average estimate ($\hat{\beta}_{j}^{\textsc{m}}$) Site-specific estimate ($\hat{\beta}_{j}$)
Variable Spatial GEE Bridge process RE Bridge process RE Gaussian process RE
Intercept 6.95 (0.39, 13.52) 2.09 (-2.74, 6.93) 2.34 (-3.06, 7.74) 2.43 (-3.07, 7.84)
Age (years) 0.22 (0.14, 0.30) 0.22 (0.14, 0.30) 0.24 (0.16, 0.33) 0.24 (0.16, 0.33)
Net-use -0.39 (-0.66, -0.11) -0.33 (-0.61, -0.05) -0.37 (-0.68, -0.06) -0.36 (-0.67, -0.06)
Treated net -0.32 (-0.65, 0.01) -0.32 (-0.67, 0.02) -0.36 (-0.75, 0.02) -0.36 (-0.75, 0.03)
Green -0.36 (-0.63, -0.08) -0.12 (-0.30, 0.07) -0.13 (-0.34, 0.07) -0.13 (-0.34, 0.08)
@@
Health center -0.26 (-0.60, 0.08) -0.26 (-0.63, 0.11) -0.29 (-0.71, 0.12) -0.30 (-0.71, 0.12)
Spatial GEE, spatial generalized estimating equation method of Cattelan and Varin, (2018); RE, random effects
The results are summarized in Table 4. Trace plots of the bridge process and Gaussian process random effect models are shown in Supplementary Figure A.5, both showing excellent convergence. The site-specific estimates using the bridge and Gaussian process random effect models give almost identical results. From the bridge process model, the empirical Bayes estimate of the attenuation factor is $\hat{\phi}=0.895$. The spatial generalized estimating equation and the proposed method also yield similar population-averaged estimates, with slight differences in magnitude for some variables. For example, using bed net is associated with approximately $32.3\%$ reduction in the population-averaged odds of malaria infection under the Spatial GEE ($e^{-0.39}=0.677$), and a 28.1% reduction in the population-averaged odds of malaria infection under the proposed bridge process random effects model ($e^{-0.33}=0.719$), holding other variables constant.
In addition to the estimated population-averaged and site-specific coefficients, we have a posterior mean $\hat{\rho}=33.3(11.2,32.5)$ and posterior mode $\hat{\rho}^{\mathrm{mode}}=10.9$ for the bridge process, and posterior mean $\hat{\rho}=29.9(13.4,40.1)$ and posterior mode $\hat{\rho}^{\mathrm{mode}}=13.5$ for the Gaussian process random effect models, where parentheses correspond to the posterior interquartile range. We assess predictive performance using the widely applicable information criterion (WAIC) (Gelman et al., , 2014) and leave-one-out cross-validation estimate using Pareto smoothed importance sampling (PSIS-LOO) (Vehtari et al., , 2024), both conditional on random effects. The WAIC is 2326.4 (standard error 39.9) for the bridge process random effect model and 2326.3 (standard error 40.0) for the Gaussian process random effect model, and the PSIS-LOO is 2326.6 (standard error 39.9) and 2326.4 (standard error 40.0) for the Gaussian process random effect model. These suggest that predictive performances are virtually identical between the proposed model and the Gaussian process random effect model, and the bridge process model remains appealing due to its dual interpretability.